OpenStudy (anonymous):

divide and simplify. assume that z is nonzero. 18/7 / (3z). the quotent is?

OpenStudy (anonymous):

18/7 / 3z = 18/7 * 1/3z = 6 / 7z or 6 (7z)^-1